the sum of two no. is 2450,if 6% one is equal to 8% of other ,the 1st no.is.
Answer by macston(5194)

x+y=2450
x=2450-y
0.06x=0.08y Substitute for x.
0.06(2450-y)=0.08y
147-0.06y=0.08y Add 0.06y to each side.
147=0.14y Divide each side by 0.14.
1050=y ANSWER 1: One of the numbers is 1050.
x=2450-y=2450-1050=1400 ANSWER 2: The other number is 1400.
CHECK:
.06(1400)=.08(1050)
84=84